# Unitarity, Causality, and Solar System Bounds, Significantly Limit Using Gravitational Waves to Test General Relativity

### v2

@@ 3.1.4 Total Amplitude
$i\mathcal{M}(\mathbf{q})=$ $\frac{i\kappa^{2}M_{1}^{2}M_{2}^{2}}{2\mathbf{q}^{2}}+\sum_{i}\Big{[}\frac{i\kappa^{4}M_{1}^{2}M_{2}^{2}}{2560\pi^{2}\tilde{\epsilon}}+\frac{i\kappa^{4}M_{1}^{2}M_{2}^{2}}{115200\pi^{2}\mathbf{q}^{6}}\biggm{(}840\kappa^{4}m_{i}^{4}\mathbf{q}^{2}+190\kappa^{4}m_{i}^{2}\mathbf{q}^{4}$
$-15\sqrt{4m_{i}^{2}\mathbf{q}^{2}+\mathbf{q}^{4}}\left(28m_{i}^{4}+4m_{i}^{2}\mathbf{q}^{2}+3\mathbf{q}^{4}\right)\log\left(\frac{\sqrt{4m_{i}^{2}\mathbf{q}^{2}+\mathbf{q}^{4}}+2m_{i}^{2}+\mathbf{q}^{2}}{2m_{i}^{2}}\right)\biggm{)}\Big{]}.$ (47)
@@ 3.2.1 Short Distance Regime
If we take the massless limit we obtain just the single term $\Delta V=-NG^{2}M_{1}M_{2}/(20\pi r^{2})$; this is given as the dashed blue curve in Figure 1. This confirms Ref. [Burns:2014bva], but is in disagreement with Ref. [Faller:2007sy] by a factor of $1/4$. We note that in Ref. [Faller:2007sy] there is no symmetry factor of $1/2$ in their bubble diagram, and they did not take into account our tadpole Feynman diagrams and the associated renormalization of the cosmological constant. This may account for the missing factor of $1/4$ from their result. For completion, we can include other contributions of massless particles, namely the massless graviton, as well as $N_{1/2}$ massless spin-$1/2$ fermions, and $N_{1}$ spin-$1$ bosons, in addition to our $N=N_{0}$ spinless bosons. This gives
@@ 3.2.2 Long Distance Regime
(where $\Delta\mathcal{M}$ means the total amplitude minus the tree level result $i\kappa^{2}M_{1}^{2}M_{2}^{2}/(2q^{2})$). We then expand the left hand side in powers of $q$ around $q=0$, and on the right hand side we perform the integral, and then expand the result in powers of $q$. However, we must truncate this at some point. We choose to expand the left hand side out to 4 terms, i.e., $\sim q^{2},\,\sim q^{4},\,\sim q^{6}$, and $\sim q^{10}$, and we truncate the expansion on the right hand side at $b_{1}$. We then obtain 4 equations for 4 constants $A,a,p,b_{1}$. The value of $b_{1}$ is not expected to be accurate at this level of truncation. However, the value of $A,a,p$ is expected to be somewhat accurate. This truncated procedure yields the values: $A\approx 0.02798$, $a\approx 1.994$, and $p\approx 2.4499$. One could try working to higher order for more precision, however, then the integrals of the terms $b_{2}/(m_{i}^{2}r^{2})$ etc. don’t convert at small $r$. So this means this procedure cannot achieve full accuracy. A more technical treatment involves the method of steepest descent, which yields the slightly altered values of [Frob:2016xte]: $A=1/(16\sqrt{\pi})$, $a=2$, and $p=2.5$. This gives the potential to leading order in the large $mr$ regime as
@@
This is given as the dashed orange curve in Figure 1. We can compare these results to those of [Burns:2014bva] where we now disagree with their result (equation (V.57)) but do agree with [Frob:2016xte].
@@ 3.3 Comparison to Solar System Bounds
where $\alpha$ is a dimensionless strength parameter and $\lambda$ is a length scale or range. Multiple experimental techniques have been used to place a bound on corrections to Newton’s potential. These are summarized in Figure 2, which is taken from Ref. [Adelberger:2003zx].
Figure 2: Experimental bounds on the size of corrections to Newton’s potential $\alpha$ as a function of scale $\lambda$; taken from Ref. [Adelberger:2003zx]. Left side focuses on microscopic scales $10^{-6}m<\lambda<10^{-2}m$, while the right side focuses on macroscopic scales $10^{-2}m<\lambda<10^{14}m$.
@@
This means the cut off on the theory obeys $1/m\gtrsim 2\times 10^{4}\,$km, which is much larger than the size of the black holes that LIGO/Virgo are studying. Hence such an effective theory could not be used in this domain. This would compromise the analysis of Ref. [Sennett:2019bpc].
@@ 4 Conclusions
Motivated by EFT cutoff bounds on higher derivative operators in gravity that was driven by gravitational wave analysis [Sennett:2019bpc], and causality bounds on the same coefficients [Caron-Huot:2022ugt], we summarized the bounds on the masses of a range of particles that are integrated out in terms of the couplings in the EFT. To determine the validity of the bound, we proceeded to compute the 1-loop correction to Newton’s potential from a massive scalar. This required renormalization of the cosmological constant and Newton’s constant which cancelled all non-local divergences. In the non-relativistic limit, this lead to a well defined amplitude whose Fourier transform gives the non-relativistic potential, including corrections to Newton’s potential. We carried out this transform numerically for all $r$ and analytically in the small $r$ and large $r$ regimes. This is summarized in Figure 1.
@@
A possible extension of our work would be to include other types of massive fields within the loop and determine their contribution to Newton’s potential. Importantly, the $\alpha_{3}$ Wilson coefficient non-zero spin vanishes [Bern:2021ppb]. However, one can then use bounds on the quartic coupling, albeit involving logarthmic corrections.
Acknowledgements
@@ A.1 Feynman Rules
Here we review the “rules and regulations” of how to perform and calculate Feynman diagrams for graviton interactions. The majority of this appendix is taken from [Choi:1994ax, Donoghue:1994dn, Donoghue:2017pgk]. As usual, the gravitational Lagrangian is expanded around a flat space background with Minkowski signature $\eta_{\mu\nu}=\text{diag}(+1,-1,-1,-1)$ given as $g_{\mu\nu}=\eta_{\mu\nu}+\kappa\,h_{\mu\nu}$, where $\kappa=\sqrt{32\pi G}$ and $h_{\mu\nu}$ is the graviton. From this expansion, along with the canonical scalar field Lagrangian, the Feynman rules can be read off. For a more in-depth procedure on their derivation please consult [Donoghue:1994dn], otherwise we simply cite the results of the Feynman rules used here.
@@ A.2 Other Equations
$F^{\mu\nu\alpha\beta}=$ $-2D\eta^{\beta\nu}q^{\mu}q^{4}q^{\alpha}-2D\eta^{\beta\mu}q^{\nu}q^{4}q^{\alpha}-2D^{3}q^{\beta}\eta^{\mu\nu}q^{4}q^{\alpha}-{8D^{2}m^{2}\eta^{\beta\nu}q^{\mu}q^{2}q^{\alpha}}-{4D^{2}q^{\beta}q^{\mu}q^{\nu}q^{2}q^{\alpha}}$
$-{8D^{2}m^{2}\eta^{\beta\mu}q^{\nu}q^{2}q^{\alpha}}-{8D^{2}m^{2}q^{\beta}\eta^{\mu\nu}q^{2}q^{\alpha}}-{48Dm^{2}q^{\beta}q^{\mu}q^{\nu}q^{\alpha}}-2D\eta^{\alpha\nu}q^{\mu}q^{4}q^{\beta}-2D\eta^{\alpha\mu}q^{\nu}q^{4}q^{\beta}$
$-{8D^{2}m^{2}\eta^{\alpha\nu}q^{\mu}q^{2}q^{\beta}}-{8D^{2}m^{2}\eta^{\alpha\mu}q^{\nu}q^{2}q^{\beta}}-2D^{3}\eta^{\alpha\beta}q^{\nu}q^{4}q^{\mu}-{8D^{2}m^{2}\eta^{\alpha\beta}q^{\nu}q^{2}q^{\mu}}+4D^{2}\eta^{\mu\nu}q^{\alpha}q^{\beta}q^{4}$
$+4D\eta^{\mu\nu}q^{\alpha}q^{\beta}q^{4}+4D^{2}\eta^{\alpha\beta}q^{\mu}q^{\nu}q^{4}+4D\eta^{\alpha\beta}q^{\mu}q^{\nu}q^{4}-16Dm^{2}\eta^{\alpha\mu}\eta^{\beta\nu}q^{4}-16Dm^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}$
$-8m^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}+16D\eta^{\beta\mu}\eta^{\alpha\nu}m^{2}q^{4}+8\eta^{\beta\mu}\eta^{\alpha\nu}m^{2}q^{4}+8\eta^{\alpha\mu}\eta^{\beta\nu}m^{2}q^{4}+16D^{2}\eta^{\alpha\beta}\eta^{\mu\nu}m^{2}q^{4}$
@@
$+2D\eta^{\beta\mu}\eta^{\alpha\nu}q^{4}q^{2}+2D\eta^{\alpha\mu}\eta^{\beta\nu}q^{4}q^{2}+2D^{3}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{2}-4D^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{2}-4D\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{2}$ (A9)
$G^{\mu\nu\alpha\beta}=$ $-8Dm^{2}\eta^{\beta\nu}q^{\mu}q^{4}q^{\alpha}-D^{3}q^{\beta}q^{\mu}q^{\nu}q^{4}q^{\alpha}-8Dm^{2}\eta^{\beta\mu}q^{\nu}q^{4}q^{\alpha}-2D^{2}q^{\beta}\eta^{\mu\nu}q^{2}q^{4}q^{\alpha}-2Dq^{\beta}\eta^{\mu\nu}q^{2}q^{4}q^{\alpha}$
@@
$-2D^{2}\eta^{\alpha\beta}q^{\nu}q^{2}q^{4}q^{\mu}-2D\eta^{\alpha\beta}q^{\nu}q^{2}q^{4}q^{\mu}+8D^{2}\eta^{\mu\nu}m^{2}q^{\alpha}q^{\beta}q^{4}-16Dm^{4}\eta^{\alpha\nu}\eta^{\beta\mu}q^{4}+2D^{2}q^{\alpha}q^{\beta}q^{\mu}q^{\nu}q^{4}$
$+8D^{2}\eta^{\alpha\beta}m^{2}q^{\mu}q^{\nu}q^{4}-16Dm^{4}\eta^{\alpha\mu}\eta^{\beta\nu}q^{4}-16Dm^{4}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}-D\eta^{\beta\mu}\eta^{\alpha\nu}q^{4}q^{4}-D\eta^{\alpha\mu}\eta^{\beta\nu}q^{4}q^{4}$
$-D^{3}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{4}+2D^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{4}+2D\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{4}+{16D\eta^{\mu\nu}m^{4}q^{\alpha}q^{\beta}q^{2}}+{16D\eta^{\beta\nu}m^{4}q^{\alpha}q^{\mu}q^{2}}$
@@
$+{16D\eta^{\alpha\beta}m^{4}q^{\mu}q^{\nu}q^{2}}+D^{3}\eta^{\mu\nu}q^{4}q^{\alpha}q^{\beta}q^{2}+D\eta^{\beta\nu}q^{4}q^{\alpha}q^{\mu}q^{2}+D\eta^{\alpha\nu}q^{4}q^{\beta}q^{\mu}q^{2}+D\eta^{\beta\mu}q^{4}q^{\alpha}q^{\nu}q^{2}$
$+D\eta^{\alpha\mu}q^{4}q^{\beta}q^{\nu}q^{2}+D^{3}\eta^{\alpha\beta}q^{4}q^{\mu}q^{\nu}q^{2}-8D^{2}m^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{2}+8D\eta^{\beta\mu}\eta^{\alpha\nu}m^{2}q^{4}q^{2}+8D\eta^{\alpha\mu}\eta^{\beta\nu}m^{2}q^{4}q^{2}$ (A10)
@@ A.3 Non-local functions $J(q^{2})$ and $B(q^{2})$
following Ref. [Donoghue:2022chi], we have defined the function that is denoted as $J(q^{2})$ and appears whenn computing the massive-bubble loop in equation (21)

### v3

@@ 3.1.4 Total Amplitude
$i\mathcal{M}(\mathbf{q})=$ $\frac{i\kappa^{2}M_{1}^{2}M_{2}^{2}}{2\mathbf{q}^{2}}+\sum_{i}\biggm{[}\frac{i\kappa^{4}M_{1}^{2}M_{2}^{2}}{2560\pi^{2}\tilde{\epsilon}}+\frac{i\kappa^{4}M_{1}^{2}M_{2}^{2}}{115200\pi^{2}\mathbf{q}^{6}}\biggm{(}840\kappa^{4}m_{i}^{4}\mathbf{q}^{2}+190\kappa^{4}m_{i}^{2}\mathbf{q}^{4}$
$-15\sqrt{4m_{i}^{2}\mathbf{q}^{2}+\mathbf{q}^{4}}\left(28m_{i}^{4}+4m_{i}^{2}\mathbf{q}^{2}+3\mathbf{q}^{4}\right)\log\left(\frac{\sqrt{4m_{i}^{2}\mathbf{q}^{2}+\mathbf{q}^{4}}+2m_{i}^{2}+\mathbf{q}^{2}}{2m_{i}^{2}}\right)\biggm{)}\biggm{]}$ (46)
@@ 3.2.1 Short Distance Regime
If we take the massless limit we obtain just the single term $\Delta V=-NG^{2}M_{1}M_{2}/(20\pi r^{2})$; this is given as the dashed blue curve in Figure 1. This confirms Ref. Burns:2014bva , but is in disagreement with Ref. Faller:2007sy by a factor of $1/4$. We note that in Ref. Faller:2007sy there is no symmetry factor of $1/2$ in their bubble diagram, and they did not take into account our tadpole Feynman diagrams and the associated renormalization of the cosmological constant. This may account for the missing factor of $1/4$ from their result. For completion, we can include other contributions of massless particles, namely the massless graviton, as well as $N_{1/2}$ massless spin-$1/2$ fermions, and $N_{1}$ spin-$1$ bosons, in addition to our $N=N_{0}$ spinless bosons. This gives
@@ 3.2.2 Long Distance Regime
(where $\Delta\mathcal{M}$ means the total amplitude minus the tree level result $i\kappa^{2}M_{1}^{2}M_{2}^{2}/(2q^{2})$). We then expand the left hand side in powers of $q$ around $q=0$, and on the right hand side we perform the integral, and then expand the result in powers of $q$. However, we must truncate this at some point. We choose to expand the left hand side out to 4 terms, i.e., $\sim q^{2},\,\sim q^{4},\,\sim q^{6}$, and $\sim q^{10}$, and we truncate the expansion on the right hand side at $b_{1}$. We then obtain 4 equations for 4 constants $A,a,p,b_{1}$. The value of $b_{1}$ is not expected to be accurate at this level of truncation. However, the value of $A,a,p$ is expected to be somewhat accurate. This truncated procedure yields the values: $A\approx 0.02798$, $a\approx 1.994$, and $p\approx 2.4499$. One could try working to higher order for more precision, however, then the integrals of the terms $b_{2}/(m_{i}^{2}r^{2})$ etc. don’t convert at small $r$. So this means this procedure cannot achieve full accuracy. A more technical treatment involves the method of steepest descent, which yields the slightly altered values of Frob:2016xte : $A=1/(16\sqrt{\pi})$, $a=2$, and $p=2.5$. This gives the potential to leading order in the large $mr$ regime as
@@
This is given as the dashed orange curve in Figure 1. We can compare these results to those of Burns:2014bva where we now disagree with their result (equation (V.57)) but do agree with Frob:2016xte .
@@ 3.3 Comparison to Solar System Bounds
where $\alpha$ is a dimensionless strength parameter and $\lambda$ is a length scale or range. Multiple experimental techniques have been used to place a bound on corrections to Newton’s potential. These are summarized in Figure 2, which is taken from Ref. Adelberger:2003zx .
Figure 2: Experimental bounds on the size of corrections to Newton’s potential $\alpha$ as a function of scale $\lambda$; taken from Ref. Adelberger:2003zx . Left side focuses on microscopic scales $10^{-6}m<\lambda<10^{-2}m$, while the right side focuses on macroscopic scales $10^{-2}m<\lambda<10^{14}m$.
@@
This means the cut off on the theory obeys $1/m\gtrsim 2\times 10^{4}\,$km, which is much larger than the size of the black holes that LIGO/Virgo are studying. Hence such an effective theory could not be used in this domain. This would compromise the analysis of Ref. Sennett:2019bpc .
@@ 4 Conclusions
Motivated by EFT cutoff bounds on higher derivative operators in gravity that was driven by gravitational wave analysis Sennett:2019bpc , and causality bounds on the same coefficients Caron-Huot:2022ugt , we summarized the bounds on the masses of a range of particles that are integrated out in terms of the couplings in the EFT. To determine the validity of the bound, we proceeded to compute the 1-loop correction to Newton’s potential from a massive scalar. This required renormalization of the cosmological constant and Newton’s constant which cancelled all non-local divergences. In the non-relativistic limit, this lead to a well defined amplitude whose Fourier transform gives the non-relativistic potential, including corrections to Newton’s potential. We carried out this transform numerically for all $r$ and analytically in the small $r$ and large $r$ regimes. This is summarized in Figure 1.
@@
A possible extension of our work would be to include other types of massive fields within the loop and determine their contribution to Newton’s potential. Importantly, the $\alpha_{3}$ Wilson coefficient non-zero spin vanishes Bern:2021ppb . However, one can then use bounds on the quartic coupling, albeit involving logarthmic corrections.
Acknowledgements.
@@ A.1 Feynman Rules
Here we review the “rules and regulations” of how to perform and calculate Feynman diagrams for graviton interactions. The majority of this appendix is taken from Choi:1994ax ; Donoghue:1994dn ; Donoghue:2017pgk . As usual, the gravitational Lagrangian is expanded around a flat space background with Minkowski signature $\eta_{\mu\nu}=\text{diag}(+1,-1,-1,-1)$ given as $g_{\mu\nu}=\eta_{\mu\nu}+\kappa\,h_{\mu\nu}$, where $\kappa=\sqrt{32\pi G}$ and $h_{\mu\nu}$ is the graviton. From this expansion, along with the canonical scalar field Lagrangian, the Feynman rules can be read off. For a more in-depth procedure on their derivation please consult Donoghue:1994dn , otherwise we simply cite the results of the Feynman rules used here.
@@ A.2 Other Equations
$F^{\mu\nu\alpha\beta}=$ $-2D\eta^{\beta\nu}q^{\mu}q^{4}q^{\alpha}-2D\eta^{\beta\mu}q^{\nu}q^{4}q^{\alpha}-2D^{3}q^{\beta}\eta^{\mu\nu}q^{4}q^{\alpha}-{8D^{2}m^{2}\eta^{\beta\nu}q^{\mu}q^{2}q^{\alpha}}$
$-{8D^{2}m^{2}\eta^{\beta\mu}q^{\nu}q^{2}q^{\alpha}}-{8D^{2}m^{2}q^{\beta}\eta^{\mu\nu}q^{2}q^{\alpha}}-{48Dm^{2}q^{\beta}q^{\mu}q^{\nu}q^{\alpha}}-2D\eta^{\alpha\nu}q^{\mu}q^{4}q^{\beta}$
$-{8D^{2}m^{2}\eta^{\alpha\nu}q^{\mu}q^{2}q^{\beta}}-{8D^{2}m^{2}\eta^{\alpha\mu}q^{\nu}q^{2}q^{\beta}}-2D^{3}\eta^{\alpha\beta}q^{\nu}q^{4}q^{\mu}-{8D^{2}m^{2}\eta^{\alpha\beta}q^{\nu}q^{2}q^{\mu}}$
$+4D\eta^{\mu\nu}q^{\alpha}q^{\beta}q^{4}+4D^{2}\eta^{\alpha\beta}q^{\mu}q^{\nu}q^{4}+4D\eta^{\alpha\beta}q^{\mu}q^{\nu}q^{4}+16D^{2}\eta^{\alpha\beta}\eta^{\mu\nu}m^{2}q^{4}$
$-8m^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}+16D\eta^{\beta\mu}\eta^{\alpha\nu}m^{2}q^{4}+8\eta^{\beta\mu}\eta^{\alpha\nu}m^{2}q^{4}+8\eta^{\alpha\mu}\eta^{\beta\nu}m^{2}q^{4}$
@@
$+2D\eta^{\beta\mu}\eta^{\alpha\nu}q^{4}q^{2}+2D\eta^{\alpha\mu}\eta^{\beta\nu}q^{4}q^{2}+2D^{3}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{2}-4D^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{2}$
$-{4D^{2}q^{\beta}q^{\mu}q^{\nu}q^{2}q^{\alpha}}-2D\eta^{\alpha\mu}q^{\nu}q^{4}q^{\beta}+4D^{2}\eta^{\mu\nu}q^{\alpha}q^{\beta}q^{4}-16Dm^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}$
$-4D\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{2}$ (A9)
$G^{\mu\nu\alpha\beta}=$ $-8Dm^{2}\eta^{\beta\nu}q^{\mu}q^{4}q^{\alpha}-D^{3}q^{\beta}q^{\mu}q^{\nu}q^{4}q^{\alpha}-8Dm^{2}\eta^{\beta\mu}q^{\nu}q^{4}q^{\alpha}-2D^{2}q^{\beta}\eta^{\mu\nu}q^{2}q^{4}q^{\alpha}$
@@
$-2D^{2}\eta^{\alpha\beta}q^{\nu}q^{2}q^{4}q^{\mu}-2D\eta^{\alpha\beta}q^{\nu}q^{2}q^{4}q^{\mu}+8D^{2}\eta^{\mu\nu}m^{2}q^{\alpha}q^{\beta}q^{4}-16Dm^{4}\eta^{\alpha\nu}\eta^{\beta\mu}q^{4}$
$+8D^{2}\eta^{\alpha\beta}m^{2}q^{\mu}q^{\nu}q^{4}-16Dm^{4}\eta^{\alpha\mu}\eta^{\beta\nu}q^{4}-16Dm^{4}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}-D\eta^{\beta\mu}\eta^{\alpha\nu}q^{4}q^{4}$
$-D^{3}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{4}+2D^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{4}+2D\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{4}+{16D\eta^{\mu\nu}m^{4}q^{\alpha}q^{\beta}q^{2}}$
@@
$+{16D\eta^{\alpha\beta}m^{4}q^{\mu}q^{\nu}q^{2}}+D^{3}\eta^{\mu\nu}q^{4}q^{\alpha}q^{\beta}q^{2}+D\eta^{\beta\nu}q^{4}q^{\alpha}q^{\mu}q^{2}+D\eta^{\alpha\nu}q^{4}q^{\beta}q^{\mu}q^{2}$
$+D\eta^{\alpha\mu}q^{4}q^{\beta}q^{\nu}q^{2}+D^{3}\eta^{\alpha\beta}q^{4}q^{\mu}q^{\nu}q^{2}-8D^{2}m^{2}\eta^{\alpha\beta}\eta^{\mu\nu}q^{4}q^{2}+8D\eta^{\beta\mu}\eta^{\alpha\nu}m^{2}q^{4}q^{2}$
$-2Dq^{\beta}\eta^{\mu\nu}q^{2}q^{4}q^{\alpha}+2D^{2}q^{\alpha}q^{\beta}q^{\mu}q^{\nu}q^{4}-D\eta^{\alpha\mu}\eta^{\beta\nu}q^{4}q^{4}+{16D\eta^{\beta\nu}m^{4}q^{\alpha}q^{\mu}q^{2}}$
$+8D\eta^{\alpha\mu}\eta^{\beta\nu}m^{2}q^{4}q^{2}+D\eta^{\beta\mu}q^{4}q^{\alpha}q^{\nu}q^{2}$ (A10)
@@ A.3 Non-local functions $J(q^{2})$ and $B(q^{2})$
following Ref. Donoghue:2022chi , we have defined the function that is denoted as $J(q^{2})$ and appears whenn computing the massive-bubble loop in equation (21)